what value of x makes this proportion true?
x/35= 3/7 I believe if i divide 35 by seven that would give the correct answer of 5. the other options are 7 11 and 15 and they dont make sense is this correct and if not how do i get the correct answer

Answer by DrBeeee(684)

Given:
(1) x/35 = 3/7
Just MULTIPLY (not divide) both sides of the equation (1) by 35 and get
(2) 35*x/35 = 35*3/7 which simplifies to
(3) x = 5*3 or
(4) x = 15
Let's check this with (1)
Is (15/35 = 3/7)?
Is (3/7 = 3/7)? Yes
Answer: the value of x that satisfies the proportion is 15.